OpenStudy (anonymous):

2(P + 1) 7 + P Solve the equation

OpenStudy (anonymous):

this is not an equation

OpenStudy (anonymous):

where is the " = "

OpenStudy (anonymous):

2(p+1)= 7+P 2p+2= 7+p -2 -2 _________________________- 2p= 5+p -p -p ___________________________- p=???

OpenStudy (anonymous):

Sorry 2(P + 1) > 7 + P

OpenStudy (anonymous):

inequality - sumthing ryt ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its the same thing , just put in the inequality symbol, where the equal sign is

OpenStudy (anonymous):

soo , it goes like this 2p+2 > 7 + p 2p+(-p) > 7 + (-2) Slove ...

OpenStudy (anonymous):

sorry * @TestTakingLifesaver

OpenStudy (anonymous):

Thanks for fanning me Tempos!